# DarkGS: Learning Neural Illumination and 3D Gaussians Relighting for Robotic Exploration in the Dark

### v1

@@ IV-C Why do we need to learn RID?
As shown in Fig. 5(a), different light has different RID patterns. Existing methods often model RID as known [25], or general functions such as the power of cosine functions [22] or a Gaussian distribution. However, for most light sources, and all light sources we used in this research, RID is not given. Modeling them with general functions such as a Gaussian distribution will not be adaptive and expressive enough to reflect the true RID of different light patterns for building a photorealistic renderer. The ablation study in Fig. 7 (columns 2 and 5) shows that while a Gaussian distribution fits the RID of flood light well, it performs much worse on the other two types of light. In comparison, our MLP-based model fits all 3 kinds of light with the best performance, showing good adaptivity to different light patterns.
@@ IV-D Why do we need to learn light falloff?
For simplicity, previous studies [24, 22, 25] use an inverse square law to model the light falloff. It is not a good fit for real-world light systems. We measure the light falloff of the center point of 3 different light sources as shown in Fig. 5(b). The plot show that when the range is close, light fall-off does not follow the inverse square law as the point light source assumption starts to fail. However, with a Lorentzian model suggested by [35], the light falloff can be better fitted with our introduced learnable parameter $\tau$. We further show numerical results in Fig. 7 (columns 2 and 3) that by learning $\tau$, the rendering performance on testing set of all 3 kinds of lights gets improved, which implies that our light fall-off model better approximates the true light falloff.
@@ IV-F Results Visualization
We deployed our system in various real-world environments and the results are shown in Fig. 8. Compared with ground truth, our model is able to reconstruct the image with photorealistic quality with learned illumination. Then we replace the light source in our model with a Lambertian light to create a global illumination that illuminates the entire FOV (FOV) of the camera with photorealistic rendering quality.
@@ V LIMITATIONS
A limitation of our work is that our model currently does not support modeling objects with heavy shadows or specular reflection (Fig. 8(a)). However, recent concurrent works [38, 11] have shown the potential to integrate them into the 3DGS framework. It is also obvious that our relighted image is in green-blue tone (Fig. 8(b)). This is because 1. We are sticking with the RAW image format throughout the pipeline, so images appear green as the RAW image nature 2. Our LED light has a high color temperature (cool color) so that the relighted image appears blue.
@@
(b) White Balancing
Figure 9: Limitations: (a) Our current method is not able to model shadows and specularity. (b) The relighted image is still in green-blue tone due to the nature of RAW images format and that white balancing is missing from our pipeline.
@@ VI CONCLUSIONS AND FUTURE WORK
This work aims to solve the problem of building 3D Gaussians and scene relighting from images taken by a moving camera-light system. Our proposed pipeline consists of NeLiS, a camera-light simulation and calibration model, and DarkGS which build’s photolistic representation for scenes in the dark. The results show that our proposed pipeline can build relightable Gaussians from images taken by the robot platform deployed in the dark, while other current approaches are not able to do the job. Ablations show that components in our proposed model effectively improve the performance of our shading model, so we can learn RID with an MLP for arbitrary light pattern, better approximate the light-falloff curve, and allow us to calibrate and deploy the system in environments that are not perfectly dark. These efforts make our model more applicable in real-world robot deployments. Future work includes modeling shadows and non-Lambertian objects, and bringing white balance and tone mapping into the loop.

### v2

@@ IV-C Why do we need to learn RID?
As shown in Fig. 6(a), different light has different RID patterns. Existing methods often model RID as known [song2020light], or general functions such as the power of cosine functions [Ma:23] or a Gaussian distribution. However, for most light sources, and all light sources we used in this research, RID is not given. Modeling them with general functions such as a Gaussian distribution will not be adaptive and expressive enough to reflect the true RID of different light patterns for building a photorealistic renderer. The ablation study in Fig. 7 (columns 2 and 5) shows that while a Gaussian distribution fits the RID of flood light well, it performs much worse on the other two types of light. In comparison, our MLP-based model fits all 3 kinds of light with the best performance, showing good adaptivity to different light patterns.
@@
Figure 9: Novel-view rendering, relighting with virtual Lambertian illumination, and manual white balance.
@@ IV-D Why do we need to learn light falloff?
For simplicity, previous studies [Ma:19, Ma:23, song2020light] use an inverse square law to model the light falloff. It is not a good fit for real-world light systems. We measure the light falloff of the center point of 3 different light sources as shown in Fig. 6(b). The plot show that when the range is close, light fall-off does not follow the inverse square law as the point light source assumption starts to fail. However, with a Lorentzian model suggested by [ryer1997light], the light falloff can be better fitted with our introduced learnable parameter $\tau$. We further show numerical results in Fig. 7 (columns 2 and 3) that by learning $\tau$, the rendering performance on testing set of all 3 kinds of lights gets improved, which implies that our light fall-off model better approximates the true light falloff.
@@ IV-F Results Visualization
We deploy our system in various real-world environments and the results are shown in Fig. 8. Compared with ground truth, our model is able to reconstruct the image with photorealistic quality with learned illumination. Then we replace the light source in our model with a virtual Lambertian light to create a global illumination that illuminates the entire FOV (FOV) of the camera with photorealistic rendering quality.
We further visualize the novel-view rendering results in Fig 9. In the first column in Fig 9, we show that the learned light effects generalize to novel views well, with visual quality comparable to training views in Fig. 8. We relight the scene (column 2) in a way similar to Fig. 8, and manual white balance (column 3).
@@ V LIMITATIONS
A limitation of our work is that our model currently does not support modeling objects with heavy shadows or specular reflection (Fig. 10(a)). However, recent concurrent works [bolanos2024gaussianshadow, R3DG2023] have shown the potential to integrate them into the 3DGS framework. It is also obvious that our relighted image is in green-blue tone (Fig. 10(b)). This is because 1. We are sticking with the RAW image format throughout the pipeline, so images appear green as the RAW image nature 2. Our LED light has a high color temperature (cool color) so that the relighted image appears blue.
@@
(b) Auto White Balancing
Figure 10: Limitations: (a) Our current method is not able to model shadows and specularity. (b) The relighted image is still in green-blue tone due to the nature of RAW images format and that automatic white balancing is missing from our pipeline.
@@ VI CONCLUSIONS AND FUTURE WORK
This work aims to solve the problem of building 3D Gaussians and scene relighting from images taken by a moving camera-light system. Our proposed pipeline consists of NeLiS, a camera-light simulation and calibration model, and DarkGS which build’s photolistic representation for scenes in the dark. The results show that our proposed pipeline can build relightable Gaussians from images taken by the robot platform deployed in the dark, while other current approaches are not able to do the job. Ablations show that components in our proposed model effectively improve the performance of our shading model, so we can learn RID with an MLP for arbitrary light pattern, better approximate the light-falloff curve, and allow us to calibrate and deploy the system in environments that are not perfectly dark. These efforts make our model more applicable in real-world robot deployments. Future work includes modeling shadows and non-Lambertian objects, and bringing automatic white balance and tone mapping into the loop.